# Overcoming Label Noise for Source-free Unsupervised Video Domain Adaptation

### v1

@@ Abstract
Despite the progress seen in classification methods, current approaches for handling videos with distribution shifts in source and target domains remain source-dependent as they require access to the source data during the adaptation stage. In this paper, we present a self-training based source-free video domain adaptation approach to address this challenge by bridging the gap between the source and the target domains. We use the source pre-trained model to generate pseudo-labels for the target domain samples, which are inevitably noisy. Thus, we treat the problem of source-free video domain adaptation as learning from noisy labels and argue that the samples with correct pseudo-labels can help us in adaptation. To this end, we leverage the cross-entropy loss as an indicator of the correctness of the pseudo-labels and use the resulting small-loss samples from the target domain for fine-tuning the model. We further enhance the adaptation performance by implementing a teacher-student framework, in which the teacher, which is updated gradually, produces reliable pseudo-labels. Meanwhile, the student undergoes fine-tuning on the target domain videos using these generated pseudo-labels to improve its performance. Extensive experimental evaluations show that our methods, termed as CleanAdapt, CleanAdapt + TS, achieve state-of-the-art results, outperforming the existing approaches on various open datasets. Our source code is publicly available at https://avijit9.github.io/CleanAdapt.
Index Terms:
@@
action recognition, domain adaptation, transfer learning, self-training
I Introduction
@@
The availability of large-scale action recognition datasets, coupled with the rise of deep neural networks, have significantly advanced the field of video understanding [1, 2, 3, 4]. Similar to other machine learning models, these action recognition models often encounter new domains with distribution-shift [5] when deployed in real-world scenarios where the data distribution of training (source domain) and test (target domain) data is different, resulting in degraded performance. A trivial solution to alleviate this problem is fine-tuning the models with labeled target domain data, which is not always feasible due to expensive target domain annotations. Unsupervised domain adaptation (UDA) tackles this problem by transferring knowledge from the labeled source domain data to the unlabeled target domain, thus eliminating the need for comprehensive annotations for the target domain [6, 7, 8, 9, 10]. Source-free UDA takes this approach one step further, where we assume that the source domain data is unavailable during the adaptation stage. This setting is more realistic than the source-dependent one primarily due to: (a) privacy concerns, it is not always possible to transfer data between the vendor (source) and the client (target), (b) storage constraints to transfer the source data to the client side (e.g., Sports-1M is about 1 TB), and (c) source-free models reduce computation time and thus cost by not using the source domain data during the adaptation stage.
@@
Figure 1: Existing approaches [11, 12, 13, 14, 15, 16, 17] have a source-dependent adaptation stage achieving marginal performance gain over the source-pretrained models. On the other hand, our proposed methods CleanAdapt and CleanAdapt + TS achieve significant performance improvements over the source-only model while being source-free (i.e., the adaptation stage does not require videos from the source domain). (Best viewed in color.)
@@
There has been a recent surge of interest in source-dependent unsupervised domain adaptation for videos [18, 11, 19, 20, 13, 12, 21, 15, 14]. These approaches either propose to directly extend the adversarial learning framework [18] from image-based methods [6] or couple it with some temporal attention weights [11, 13] and self-supervised pretext tasks [13, 12] to align the segment-level features between the domains. However, these strategies produce only a modest $\sim 2\%$ gain over the source-only model (see Figure 1). Recently, there has been a paradigm shift from adversarial to contrastive learning framework [15, 14, 16] for video domain adaptation. As shown in Figure 1, CoMix [16] achieves 6.4% and 5.1% gain over the source-only model on the UCF $\rightarrow$ HMDB and HMDB $\rightarrow$ UCF datasets, respectively. However, all these methods are inherently complex and use source domain videos during the adaptation stage, which is untenable in several scenarios [22, 23, 24], as discussed earlier. Due to its practical relevance, source-free domain adaptation is a well-known problem for different computer vision tasks such as image classification [25, 26], and semantic segmentation [27] but relatively under-explored in the context of the video classification task. Therefore, there is a need to investigate source-free domain adaptation for video classification tasks in order to improve the practicality and efficiency of today’s approaches.
Figure 2: Performance of our proposed methods CleanAdapt and CleanAdapt + TS as compared to source-only and target supervised models. (Best viewed in color.)
@@
In this work, we present an effective approach that leverages the self-training framework [28] for source-free video UDA where we do not have access to the source-domain videos during the adaptation stage. We generate pseudo-labels for the unlabeled target domain videos using a source pre-trained model. These pseudo-labels are indeed noisy due to the existing domain gap. Finetuning the source pre-trained model with these noisy pseudo-labels is a sub-optimal solution as the presence of incorrect pseudo-labels hinders the adaptation stage, as discussed in Sec. IV-D. However, we observe that these pseudo-labeled target domain videos are not completely unusable, and in fact, there is a substantial number of target domain videos with correct pseudo-labels. For example, in the case of HMDB $\rightarrow$ UCF, the HMDB pre-trained model produces pseudo-labels with $\sim 90\%$ accuracy on the UCF dataset, and we experimentally show that this amount of data is sufficient for adaptation. Throughout this paper, we refer to these samples with correct pseudo-labels as clean, whereas the samples with incorrect pseudo-labels are termed noisy. We observe that the underlying network learns clean samples first before memorizing the noisy samples, and this acts as the core idea behind the adaptation stage in our proposed method (Figure 3). We discuss this in Sec. III-C.
@@
We treat this problem as learning from noisy labels and propose a self-training based approach that selects clean samples from the noisy pseudo-labeled target domain samples to re-train the model for gradually adapting to the target domain in an iterative manner. This method progressively adapts the model to the target domain in an iterative fashion. Thus, we name our approach as CleanAdapt. Additionally, we employ a teacher-student network [29, 30] to produce more resilient pseudo-labels, where the teacher network is continuously updated by incorporating the temporal ensemble of student networks. This approach generates more consistent pseudo-labels, thus aiding the enhancement of the student network. We refer to this version of our method as CleanAdapt + TS. Our proposed methods surpass all other source-dependent state-of-the-art methods by a large margin on UCF $\leftrightarrow$ HMDB and EPIC-Kitchens datasets, despite being source-free (see Figure 1 and Figure 2).
@@
An earlier version of our approach was published as a conference paper [31]. This manuscript proposes the following significant improvements.
@@
    We comprehensively survey existing video domain adaptation approaches for different domain adaptation setups in Sec. II.
@@
    In Sec. III, we propose an improved version of our pseudo-label generation process using a teacher-student framework. This modification yields reliable pseudo-label generation, thereby enhancing the overall performance.
@@
    We conduct a comprehensive analysis of the results obtained with the popular UCF $\leftrightarrow$ HMDB and EPIC-Kitchens datasets and detailed in Sec. IV.
@@
    We also include evaluations and comparisons in Sec. IV with recently developed source-free video domain adaptation techniques like [32, 17, 33]. Our CleanAdapt and CleanAdapt + TS outperform all the existing video domain adaptation approaches.
The paper is organized as follows. In Section II, we review some of the notable and recent works in this domain. Section III describes the CleanAdapt and CleanAdapt + TS methods. In Section IV, we present the experimental analyses and finally conclude in Section V.
II Related Work
@@
Convolutional neural networks (CNNs) are now the de facto solution for action recognition tasks. Various efforts have been made in this context to capture spatio-temporal information in videos, starting from two-stream networks with 2D [1, 34, 2, 35, 36, 37, 38, 39] to 3D CNNs [3, 40, 35, 41, 42, 43, 44]. Recent advances in action recognition focus on capturing long-term context from videos [45, 46, 47, 48, 49, 50]. However, a common limitation of existing methods is their dependence on training data that closely matches the distribution of the test data. When there is a subtle difference in the distribution between the training and the test domains, these models struggle to generalize effectively. Consequently, fine-tuning with a large amount of labeled data from the target domain is often required, which can be both time-consuming and expensive. To address this issue, our focus is on unsupervised video domain adaptation, aiming to overcome the need for labeled target domain data.
TABLE I: Summary of domain adaptation for action recognition methods. “N/A” denotes the unavailability of source code.
Serial No. Year Methods Venue Code Link Backbone Core Components Types of Videos
@@
1 2018 AMLS  [18] BMVC N/A C3D Subspace alignment Third person
2 2018 DAAAA  [18] BMVC N/A C3D Domain invariant feature learning; Adversarial learning with domain discriminator Third person
3 2019 TA3N  [11] ICCV PyTorch ResNet-101 Attention alignment; Temporal discrepancy; Adversarial alignment Third person
4 2020 TCoN  [20] AAAI N/A BN-Inception, C3D Cross-domain co-attention; Adversarial learning for temporal adaptation Third person
5 2020 MM-SADA  [12] CVPR Tensorflow I3D Self-supervised cross-modal alignment; Adversarial Learning First person
6 2020 Choi et al. [19] WACV N/A I3D Adversarial learning Third-person
7 2020 SAVA  [13] ECCV N/A I3D Align important clips; Self-supervised clip-order prediction Third-person
8 2021 STCDA  [14] CVPR N/A BN-Inception, I3D Spatio-temporal contrastive learning; pseudo-labeling First and Third person
9 2021 Kim et al. [51] ICCV N/A I3D Contrastive learning; cross-modal and cross-domain alignment First and Third person
10 2021 CoMix  [16] NeurIPS PyTorch I3D Temporal contrastive learning; background mixing; pseudo-labeling First and Third person
11 2021 MAN  [52] TC N/A ResNet-152 Multiple-view adversarial learning network Third person
12 2022 CO2A  [21] WACV PyTorch I3D Dual head contrastive network Third person and Synthetic
@@
14 2022 CIA  [54] CVPR N/A I3D Cross-modal complementarity and consensus First and Third person
15 2022 ACAN  [55] TNNLS N/A I3D, MFNet Pixel correlation alignment Third person
16 2022 TranSVAE  [56] Arxiv PyTorch I3D Disentanglement of domain-related and semantic-related information First and Third person
17 2022 Wu et al. [57] Neurocomputing N/A ResNet-50 Video-level mix-up learning; Dynamic sampling strategy Third person
18 2022 Zhang et al. [58] CVPR PyTorch ConvLSTM, I3D, Timesformer Frame dependency modeling Third person
19 2022 MD-DMD et al. [17] MM N/A I3D Dynamic modal distillation First and Third person
20 2023 Broome et al. [59] WACV N/A SlowFast (video), Resnet-18 (audio) Audio-adaptive encoder First and Third person
@@
21 2022 CleanAdapt [31] ICVGIP N/A I3D Pseudo-labeling; Learning from noisy labels First and Third person
22 2022 ATCoN [32] ECCV PyTorch ResNet-50 Temporal consistency network Third person
23 2022 MTRAN  [33] MM N/A I3D, Transformer Temporal relative alignment; Mix-up First and Third person
Partial Video Domain Adaptation
24 2021 PTAN [60] ICCV PyTorch Resnet-50 Temporal attentive network; Adversarial learning Third person
25 2022 MCAN [61] MM N/A TSN Cluster-calibrated partial adversarial network Third person
Open-Set Video Domain Adaptation
26 2021 Chen et al. [62] MM-Asia PyTorch Resnet-101 Class-conditional extreme value theory Third person
27 2021 DMD [63] ICASSP N/A Resnet-101 Dual-metric discriminator Third person
Multi-Source Video Domain Adaptation
28 2021 TAMAN [64] Arxiv N/A TRN Temporal attentive moment alignment Third person
Test-time Video Domain Adaptation
29 2022 Azimi et al. [65] WACV N/A - Test-time adaptation; Self-supervised learning Third person
30 2023 ViTTA  [66] CVPR PyTorch ResNet50, Video-Swin Estimation of test set statistics towards the training statistics Third person
@@
Early works [18, 11, 20, 12, 13] on video UDA are inspired by the adversarial framework [6] for image-based UDA tasks. Jamal et al. [18] proposes to align the source and the target domains using a subspace alignment technique and outperform all the previous shallow methods. Chen et al. [11] show the efficacy of attending to the temporal dynamics of video for domain adaptation. TCoN [20] is a cross-domain co-attention module for matching the source and the target domain features with appearance and motion streams. Munro et al. [12] were among the first to show the effectiveness of learning multi-modal correspondence for video domain adaptation. SAVA [13] is an attention-augmented model with a clip order prediction task to re-validate the effectiveness of self-supervised learning for video domain adaptation, as shown in [12]. Overall, the adversarial methods are complex and sensitive to the choice of hyperparameters [16].
@@
Figure 3: Average cross-entropy loss per epoch of training with pseudo-labeled target domain videos for clean vs. noisy samples with (a) RGB modality and (b) Flow modality. We term the target domain samples with correct pseudo-labels as clean samples and with incorrect pseudo-labels as noisy samples. Note that, the groundtruth labels are only used to identify the clean vs. the noisy samples for visualization purposes and not used for training the model. Deep neural networks learn the clean samples first before memorizing the noisy samples according to the deep memorization effect as proposed in [67]. In our proposed approach CleanAdapt, we exploit this connection to select the clean samples for fine-tuning the model to adapt to the target domain. (Best viewed in color.)
@@
There has been a recent shift from adversarial to contrastive learning-based methods for the video UDA task. Song et al. [14] propose to bridge the domain gap using a self-supervised contrastive framework named cross-modal alignment. In a similar direction, Kim et al. [15] use a cross-modal feature alignment loss for learning a domain adaptive feature representation. CoMix [16] represents videos as graphs and uses temporal-contrastive learning over graph representations for transferable feature learning. Additionally, these methods [14, 15, 16] generate pseudo-labels from the source pre-trained model for the target domain videos and use only the target domain samples with high-confident pseudo-labels in their contrastive loss in each iteration. However, the source-only model often makes incorrect predictions with high confidence due to the distribution shift for target domain videos, which can hinder adaptation. To address this, we treat target pseudo-labels as noisy and formulate the domain adaptation problem as learning from noisy labels. Moreover, the adaptation stage in these methods [18, 11, 19, 20, 13, 12, 14, 15, 16] is source-dependent. This is an impractical requirement as source data transfer during the deployment phase of the model is often infeasible.
@@
Recently, ATCoN [32] and our conference paper CleanAdapt [31] have addressed this issue of source data dependency. These methods introduce a source-free adaptation approach, i.e., it does not rely on source domain videos during the adaptation stage. Other works  [65, 64, 63, 62, 61, 60] have made notable advancements in tackling various domain adaptation scenarios for videos, including open-set, partial-set, and test-time setups. In Table I, we provide an overview of existing methods for video domain adaptation.
@@
Self-training based methods with careful design choices may still produce over-confident incorrect predictions. To alleviate this issue, we resort to learning from label-noise literature. One of the popular approaches to reducing the effect of noisy-labels is to design noise-robust losses [68, 69, 70]. However, these methods fail to handle real-world noise [71]. According to [67], deep neural networks produce small loss values for samples with correct pseudo-labels. Thus, a popular direction for handling label-noise is to use the cross-entropy loss to indicate label correctness [72, 73] and leverage these small-loss samples for re-training the networks. In this work, we demonstrate that the small-loss samples are potentially clean samples and are effective in helping our source pre-trained model adapt to the target domain if these samples are used for fine-tuning. Therefore, our proposed approach is simpler that the existing approaches, requiring solely pseudo-labeled samples from the target domain.
III Approach
III-A Problem Definition

### v2

# Source-free Video Domain Adaptation by Learning from Noisy Labels

@@ Abstract
Despite the progress seen in classification methods, current approaches for handling videos with distribution shifts in source and target domains remain source-dependent as they require access to the source data during the adaptation stage. In this paper, we present a self-training based source-free video domain adaptation approach to address this challenge by bridging the gap between the source and the target domains. We use the source pre-trained model to generate pseudo-labels for the target domain samples, which are inevitably noisy. Thus, we treat the problem of source-free video domain adaptation as learning from noisy labels and argue that the samples with correct pseudo-labels can help us in adaptation. To this end, we leverage the cross-entropy loss as an indicator of the correctness of the pseudo-labels and use the resulting small-loss samples from the target domain for fine-tuning the model. We further enhance the adaptation performance by implementing a teacher-student (TS) framework, in which the teacher, which is updated gradually, produces reliable pseudo-labels. Meanwhile, the student undergoes fine-tuning on the target domain videos using these generated pseudo-labels to improve its performance. Extensive experimental evaluations show that our methods, termed as CleanAdapt, CleanAdapt + TS, achieve state-of-the-art results, outperforming the existing approaches on various open datasets. Our source code is publicly available at https://avijit9.github.io/CleanAdapt.
keywords:
@@
action recognition, domain adaptation, transfer learning, and self-training.
1 Introduction
@@
The availability of large-scale action recognition datasets, coupled with the rise of deep neural networks, have significantly advanced the field of video understanding [carreira2017quo]. Similar to other machine learning models, these action recognition models often encounter new domains with distribution-shift when deployed in real-world scenarios where the data distribution of training (source domain) and test (target domain) data is different, resulting in degraded performance. A trivial solution to alleviate this problem is fine-tuning the models with labeled target domain data, which is not always feasible due to expensive target domain annotations. Unsupervised domain adaptation (UDA) tackles this problem by transferring knowledge from the labeled source domain data to the unlabeled target domain, thus eliminating the need for comprehensive annotations for the target domain [ganin2015unsupervised]. Source-free UDA takes this approach one step further, where we assume that the source domain data is unavailable during the adaptation stage. This setting is more realistic than the source-dependent one primarily due to (a) privacy concerns, it is not always possible to transfer data between the vendor (source) and the client (target), (b) storage constraints to transfer the source data to the client side (e.g., Sports-1M is about 1 TB), and (c) source-free models reduce computation time and thus cost by not using the source domain data during the adaptation stage.
@@
Figure 1: Existing approaches have a source-dependent adaptation stage achieving marginal performance gain over the source-pretrained models. On the other hand, our proposed methods CleanAdapt and CleanAdapt + TS achieve significant performance improvements over the source-only model while being source-free (i.e., the adaptation stage does not require videos from the source domain). (Best viewed in color.)
@@
There has been a recent surge of interest in source-dependent unsupervised domain adaptation for videos [kim2021learning, song2021spatio, 10139790]. These approaches either propose to directly extend the adversarial learning framework [jamal2018deep] from image-based methods [ganin2015unsupervised] or couple it with some temporal attention weights [chen2019temporal, choi2020shuffle] and self-supervised pretext tasks [choi2020shuffle, munro2020multi] to align the segment-level features between the domains. However, these strategies produce only a modest $\sim 2\%$ gain over the source-only model (see Figure 1). Recently, there has been a paradigm shift from adversarial to contrastive learning framework [kim2021learning, song2021spatio, sahoo2021contrast] for video domain adaptation. As shown in Figure 1, CoMix [sahoo2021contrast] achieves 6.4% and 5.1% gain over the source-only model on the UCF $\rightarrow$ HMDB and HMDB $\rightarrow$ UCF datasets, respectively. However, all these methods are inherently complex and use source domain videos during the adaptation stage, which is untenable in several scenarios [yang2021exploiting], as discussed earlier. Due to its practical relevance, source-free domain adaptation is a well-known problem for different computer vision tasks such as image classification [yang2021exploiting], and semantic segmentation [guan2021scale] but relatively under-explored in the context of the video classification task. Therefore, there is a need to investigate source-free domain adaptation for video classification tasks in order to improve the practicality and efficiency of today’s approaches.
Figure 2: The radar plot illustrates the performance improvements of our proposed methods, CleanAdapt and CleanAdapt + TS (shown in orange and red, respectively), compared to the source-only model on multiple benchmarks. The source-only model (shown in yellow), trained on the source domain and tested on the target domain, serves as the lower bound of adaptation performance, while the target-supervised model (shown in blue), trained and tested on target domain videos, represents the upper bound. (Best viewed in color.)
@@
In this work, we present an effective approach that leverages the self-training framework for source-free video UDA, where we do not have access to source-domain videos during the adaptation stage. We generate pseudo-labels for the unlabeled target domain videos using a source pre-trained model. These pseudo-labels are indeed noisy due to the existing domain gap. Finetuning the source pre-trained model with these noisy pseudo-labels is a sub-optimal solution as the presence of incorrect pseudo-labels hinders the adaptation stage, as discussed in Sec. 4.4. However, we observe that these pseudo-labeled target domain videos are not completely unusable, and in fact, there is a substantial number of target domain videos with correct pseudo-labels. For example, in the case of HMDB $\rightarrow$ UCF, the HMDB pre-trained model produces pseudo-labels with $\sim 90\%$ accuracy on the UCF dataset, and we experimentally show that this amount of data is sufficient for adaptation. Throughout this paper, we refer to these samples with correct pseudo-labels as clean, whereas the samples with incorrect pseudo-labels are termed noisy.
@@
We treat the problem of source-free domain adaptation as learning from noisy labels and propose a self-training based approach that selects clean samples from the noisy pseudo-labeled target domain samples to re-train the model for gradually adapting to the target domain in an iterative manner. We observe that deep neural networks tend to learn from the clean samples first before memorizing the noisy samples in the later stage of training according to the deep memorization effect [arpit2017closer]. In Figure 3, we validate this effect empirically for both appearance (RGB) and motion (flow) modalities. From Figure 3, it is clear that the model produces small-loss for the clean samples. In contrast to that, the noisy samples have high loss values for both the modalities. In our paper, we exploit this connection between small-loss and clean instances and propose an approach for source-free video domain adaptation. We discuss this in detail in Sec. 3.3. This method progressively adapts the model to the target domain in an iterative fashion. Thus, we name our approach as CleanAdapt. Additionally, we employ a teacher-student network [sohn2020fixmatch] to produce more resilient pseudo-labels, where the teacher network is continuously updated by incorporating the temporal ensemble of student networks. This approach generates more consistent pseudo-labels, thus aiding the enhancement of the student network. We refer to this version of our method as CleanAdapt + TS. Our proposed methods surpass all other source-dependent state-of-the-art methods by a large margin on UCF $\leftrightarrow$ HMDB and EPIC-Kitchens datasets, despite being source-free (see Figure 1 and Figure 2).
@@
An earlier version of our approach was published as a conference paper [dasgupta2022overcoming]. This manuscript proposes the following significant improvements.
@@
    We comprehensively survey existing video domain adaptation approaches for different domain adaptation setups in Sec. 2.
@@
    We propose a strategy to filter out the incorrect pseudo-labeled (noisy) samples based on the deep memorization effect [arpit2017closer] and utilize the clean samples in the adaptation stage.
@@
    In Sec. 3, we propose an improved version of our pseudo-label generation process using a teacher-student framework. This modification yields reliable pseudo-label generation, thereby enhancing the overall performance.
@@
    We conduct a comprehensive analysis of the results obtained with the popular UCF $\leftrightarrow$ HMDB and EPIC-Kitchens datasets and detailed in Sec. 4.
  5. 5.
    We also include evaluations and comparisons in Sec. 4 with recently developed source-free video domain adaptation techniques like [xu2022source, yin2022mix, huang2022relative]. Our CleanAdapt and CleanAdapt + TS outperform all the existing video domain adaptation approaches.
The paper is organized as follows. In Section 2, we review some of the notable and recent works in this domain. Section 3 describes the CleanAdapt and CleanAdapt + TS methods. In Section 4, we present the experimental analyses and finally conclude in Section 5.
2 Related Work
@@
Convolutional neural networks (CNNs) are now the de facto solution for action recognition tasks. Various efforts have been made in this context to capture spatio-temporal information in videos, starting from two-stream networks with 2D [zhou2018temporal] to 3D CNNs [carreira2017quo]. However, a common limitation of existing methods is their dependence on training data that closely matches the distribution of the test data. When there is a subtle difference in the distribution between the training and the test domains, these models struggle to generalize effectively. Consequently, fine-tuning with a large amount of labeled data from the target domain is often required, which can be both time-consuming and expensive. To address this issue, our focus is on unsupervised video domain adaptation, aiming to overcome the need for labeled target domain data.
Table 1: Summary of domain adaptation for action recognition methods. “N/A” denotes the unavailability of source code.
Methods Venue Code Link Backbone Core Components Types of Videos
@@
AMLS  [jamal2018deep] BMVC’18 N/A C3D Subspace alignment Third person
DAAAA  [jamal2018deep] BMVC’18 N/A C3D Domain invariant feature learning; Adversarial learning with domain discriminator Third person
TA3N  [chen2019temporal] ICCV’19 PyTorch ResNet-101 Attention alignment; Temporal discrepancy; Adversarial alignment Third person
TCoN  [pan2020adversarial] AAAI’20 N/A BN-Inception, C3D Cross-domain co-attention; Adversarial learning for temporal adaptation Third person
$M$-SADA  [munro2020multi] CVPR’20 Tensorflow I3D Self-supervised cross-modal alignment; Adversarial Learning First person
Choi et al[choi2020unsupervised] WACV’20 N/A I3D Adversarial learning Third-person
SAVA  [choi2020shuffle] ECCV’20 N/A I3D Align important clips; Self-supervised clip-order prediction Third-person
STCDA  [song2021spatio] CVPR’21 N/A BN-Inception, I3D Spatio-temporal contrastive learning; pseudo-labeling First and Third person
Kim et al[kim2021domain] ICCV’21 N/A I3D Contrastive learning; cross-modal and cross-domain alignment First and Third person
CoMix  [sahoo2021contrast] NeurIPS’21 PyTorch I3D Temporal contrastive learning; background mixing; pseudo-labeling First and Third person
CO2A  [da2022dual] WACV’22 PyTorch I3D Dual head contrastive network; Synthetic data Third person
@@
CIA  [yang2022interact] CVPR’22 N/A I3D Cross-modal complementarity and consensus First and Third person
TranSVAE  [wei2022unsupervised] NeurIPS’23 PyTorch I3D Disentanglement of domain-related and semantic-related information First and Third person
MD-DMD et al[yin2022mix] MM’22 N/A I3D Dynamic modal distillation First and Third person
Broome et al[broome2023recur] WACV’23 N/A SlowFast (video), Resnet-18 (audio) Audio-adaptive encoder First and Third person
CTAN [10139790] TCSVT’23 PyTorch I3D Channel-temporal attention network First person
@@
CleanAdapt [dasgupta2022overcoming] ICVGIP’22 PyTorch I3D Pseudo-labeling; Learning from noisy labels First and Third person
ATCoN [xu2022source] ECCV’22 PyTorch ResNet-50 Temporal consistency network Third person
MTRAN  [huang2022relative] MM’22 N/A I3D, Transformer Temporal relative alignment; Mix-up First and Third person
@@
Early works [chen2019temporal, munro2020multi, choi2020shuffle] on video UDA are inspired by the adversarial framework [ganin2015unsupervised] for image-based UDA tasks. Jamal et al[jamal2018deep] proposes to align the source and the target domains using a subspace alignment technique and outperform all the previous shallow methods. Chen et al[chen2019temporal] show the efficacy of attending to the temporal dynamics of video for domain adaptation. TCoN [pan2020adversarial] is a cross-domain co-attention module for matching the source and the target domain features with appearance and motion streams. Munro et al[munro2020multi] were among the first to show the effectiveness of learning multi-modal correspondence for video domain adaptation. SAVA [choi2020shuffle] is an attention-augmented model with a clip order prediction task to re-validate the effectiveness of self-supervised learning for video domain adaptation, as shown in [munro2020multi]. Overall, the adversarial methods are complex and sensitive to the choice of hyperparameters [sahoo2021contrast].
@@
Figure 3: Average cross-entropy loss per epoch of training with pseudo-labeled target domain videos for clean vs. noisy samples with (a) RGB modality and (b) Flow modality. We term the target domain samples with correct pseudo-labels as clean samples and with incorrect pseudo-labels as noisy samples. Note that, the groundtruth labels are only used to identify the clean vs. the noisy samples for visualization purposes and not used for training the model. Deep neural networks learn the clean samples first before memorizing the noisy samples according to the deep memorization effect as proposed in [arpit2017closer]. In our proposed approach CleanAdapt, we exploit this connection to select the clean samples for fine-tuning the model to adapt to the target domain. (Best viewed in color.)
@@
There has been a recent shift from adversarial to contrastive learning-based methods for the video UDA task. Song et al[song2021spatio] propose to bridge the domain gap using a self-supervised contrastive framework named cross-modal alignment. In a similar direction, Kim et al[kim2021learning] use a cross-modal feature alignment loss for learning a domain adaptive feature representation. CoMix [sahoo2021contrast] represents videos as graphs and uses temporal-contrastive learning over graph representations for transferable feature learning. Additionally, these methods [song2021spatio, kim2021learning, sahoo2021contrast] generate pseudo-labels from the source pre-trained model for the target domain videos and use only the target domain samples with high-confident pseudo-labels in their contrastive loss in each iteration. However, the source-only model often makes incorrect predictions with high confidence due to the distribution shift for target domain videos, which can hinder adaptation. To address this, we treat target pseudo-labels as noisy and formulate the domain adaptation problem as learning from noisy labels. Moreover, the adaptation stage in these methods [song2021spatio, kim2021learning, sahoo2021contrast] is source-dependent. This is an impractical requirement as source data transfer during the deployment phase of the model is often infeasible.
@@
Recently, ATCoN [xu2022source] and our conference paper CleanAdapt [dasgupta2022overcoming] have addressed this issue of source data dependency. These methods introduce a source-free adaptation approach, i.e., it does not rely on source domain videos during the adaptation stage. In Table 1, we provide an overview of existing methods for unsupervised and source-free video domain adaptation. Xu et al. [xu2022video] provide an extensive survey of video domain adaptation, encompassing a variety of setups.
@@
Self-training based methods with careful design choices may still produce over-confident, incorrect predictions. To alleviate this issue, we resort to learning from label-noise literature. One of the popular approaches to reducing the effect of noisy-labels is to design noise-robust losses [feng2021can]. However, these methods fail to handle real-world noise [zhang2021prototypical]. According to [arpit2017closer], deep neural networks produce small loss values for samples with correct pseudo-labels. Thus, a popular direction for handling label-noise is to use the cross-entropy loss to indicate label correctness [han2018co] and leverage these small-loss samples for re-training the networks. In this work, we demonstrate that the small-loss samples are potentially clean samples and are effective in helping our source pre-trained model adapt to the target domain if these samples are used for fine-tuning. Therefore, our proposed approach is simpler that the existing approaches, requiring solely pseudo-labeled samples from the target domain.
3 Approach
Figure 4: Overview of the three stages of our CleanAdapt + TS framework for source-free video domain adaptation, which has three stages. (a) The model ($f_{a}$) is first pre-trained on the labeled source domain videos from $\mathcal{D}_{s}$. For brevity, only the single-stream model is shown here. (b) This source pre-trained model is then used to generate pseudo-labels $\hat{y}$ for the unlabeled target domain videos from $\mathcal{D}_{t}$. Inevitably, these pseudo-labels are noisy due to the domain shift between the source and the target domains. (c) A clean sample selection module is used to select a set $\mathcal{D}_{cl}$ of small-loss samples as potential clean samples. The source pre-trained model is finetuned on these clean samples from $\mathcal{D}_{cl}$ using their corresponding pseudo-labels $\hat{y}$. We repeat this step multiple times. See Sec. 4 for implementation details. (Best viewed in color.)
3.1 Problem Definition